part 1: a spectrometer uses a grating with 11600 lines/ cm. Find the angle at which red light, 632.3 nm, has the first-order bright band. Answer in units of degrees.

part 2: find the angle at which blue light, 419 nm, has the first-order bright band. Answer in units of degrees.
Answer #1

S = 1 x 10^7/11600 = 862.07 nm

A)

Theta_red = arcsin (632.3/862.07) = 47.18 degree

B)

Theta_blue = arcsin (419/862.07) = 29.08 degree

Comment in case any doubt please rate my answer....